A point moves along an arc of a circle of radius R. Its velocity depends on the distance covered s as v = a sqrt(s), where a is a constant. Find the angle α between the vector of the total acceleration and the vector of velocity as a function of s.
1
Expert's answer
2020-10-12T07:49:32-0400

Since "V=A\times \sqrt S"

"a_r=\frac {V^2}{R}=\frac {A^2\times S}{R}" ,

"a_t=\frac {dV}{dt}=\frac{A}{2\times \sqrt S}" "\times \frac{dS}{dt}=\frac {A}{2\times \sqrt S}" "\times A \sqrt S= \frac {A^2}{2}" .

Since "a_t" is a positive constant, the particle velocity increases with time, and the tangential acceleration vector and the velocity vector coincide in direction. Therefore, the angle between "\vec V" and "\vec a" (total acceleration) is equal to, the angle between "a_t\times \hat v_t" and "\vec a" , and it can be found "\u03b1" by the formula: "\tg {\alpha}=\frac {\lvert{a_n}\lvert}{\lvert a_t \lvert}" "=\frac {\frac {a^2\times s}{R}}{0.5\times a^2}""=\frac {2\times S}{R}" , therefore "\alpha=\arctg \frac {2\times s}{R}" .